Solution (source code)

= Solution

Conditioning on $U=X+Y$, <independence> of $Z$ gives $\mathbb P(Z>u)=1-u$ for $0\le u\le1$ and zero for $u>1$. Using the preceding <probability density function>,
$$
\boxed{\mathbb P(Z>X+Y)=\int_0^1(1-u)u\,du=\frac12-\frac13=\frac16.}
$$